Cost of Equity: Dividend Growth

Summerdahl Resort's common stock is currently trading at $28.00 a share. The stock is expected to pay a dividend of $1.75 a share at the end of the year (D1 = $1.75), and the dividend is expected to grow at a constant rate of 8% a year. What is the cost of common equity? Round your answer to two decimal places.
Answer #1

Current market price (P0) = $28

Dividend at the end of the year (D1) = $1.75

Growth rate (g) = 0.08

Cost of common equity (Ke) = (D1/P0) + g

= (1.75/28) + 0.08

= 0.0625 + 0.08

= 0.1425

= 14.25%

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.